OpenStudy (anonymous):

Assume f(x)= 2x+8 and g(x)= 5. What is the value of (f+g)(3)?

OpenStudy (aum):

(f+g)(3) = f(3) + g(3) f(x) = 2x + 8 f(3) = ? g(x) = 5 g(3) = ?

OpenStudy (anonymous):

is that the answer just 5?